Suppose milk and honey are the only product produced in Ghana. Use the data below to answer the questions below.
2011, milk price GH@1 and quantity 100, honey price GH@2 and quantity 50. 2012, milk price gh@1 and quantity 200, honey price gh@2 and quantiy 100 .2013 , milk price gh@2 and quantity 200, honey price GH@4 and quantity 50
Base year; 2011 required
Compute for each year the nominal,real deflator of the GDP.
Compute for 2012 and 2013 from the respective proceeding years the percentage change in nominal,real and deflator of the GDP.
1
Expert's answer
2021-08-02T15:15:53-0400

Real GDP

2011: "(1\times 100)+(2\times 50)=200"

2012: "(1\times 200)+(2\times 100)=400"

2013: "(2\times 200)+ (4\times 50)=600"

Nominal GDP

2011: "(1\times 100)+(2\times50)=200"

2012: "(1\times 200)+ (2\times 100)=400"

2013: "(1\times 200)+ (2\times50)=300"

Deflator

"=\frac {Nominal GDP}{Real GDP}\times 100"


2011:"=\frac {200}{200}\times100=100" %

2012: "=\frac {400}{400}\times100=100"%

2013: "=\frac {300}{600}\times 100=50" %

% change in nominal GDP

2011-2012: "=\frac {400-200}{200}\times100=100"%

2012-2013: "=\frac {300-400}{400}\times100=-25"%

% change in real GDP

2011-2012: "=\frac {400-200}{200}\times100=100"%

2012-2013: "=\frac {600-400}{400}\times100=50"%

% change in deflator

2011-2012: "=100-100=0"%

2012-2013: "=50-100=-50"%